OpenStudy (ksaimouli):

simplify 1/1-sinx +1/sinx using basic idenites

OpenStudy (anonymous):

so it looks like this: \[\frac{1}{1- sin(x)} + \frac{1}{sin(x)} \]

OpenStudy (ksaimouli):

no

OpenStudy (ksaimouli):

the second one in denominayor 1+sinx

OpenStudy (anonymous):

\[\frac{1}{1-sin(x)} + \frac{1}{sin(x)+1} \]

OpenStudy (anonymous):

ok now tell me what identity would you use to cancel out the denominator?

OpenStudy (ksaimouli):

+1 should be left side

OpenStudy (ksaimouli):

1+sinx

OpenStudy (ksaimouli):

not sinx +1

OpenStudy (anonymous):

\[\frac{1}{1+sin(x)} + \frac{1}{sin(x)} \]

OpenStudy (ksaimouli):